OpenStudy (buttermequeasy):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30). 10 0 5 −10

OpenStudy (anonymous):

10

OpenStudy (anonymous):

30/3=m

OpenStudy (anonymous):

Its (y2-y1)/(x2-x1)=(30-0)/(3-0)=30/3=10

OpenStudy (buttermequeasy):

Thank you.

OpenStudy (anonymous):

No prob

OpenStudy (anonymous):

well done!